# NoiseSDF2NoiseSDF: Learning Clean Neural Fields from Noisy Supervision

### v1

@@ 5.3 Results on ShapeNet
We compared our method against 3DS2V [48] and 3DILG [47] on the seven largest ShapeNet subsets, using the same training and testing data split. We used the official pretrained models released by the respective authors. Evaluation results are reported for each subset at noise levels of 0.01 (Table 1) and 0.02 (Table 2); see visualization results in Figure 3. The results demonstrate that our method maintains robustness under both mild $\sigma=0.01$ and severe $\sigma=0.02$ corruption levels. Under lower corruption ($\sigma=0.01$), our method outperforms all competing methods across all evaluation metrics. For instance, IoU rises by 4% for chair and by about 9% for rifle. Chamfer Distance drops from 0.010 to 0.008 for airplane and from 0.011 to 0.009 for lamp. Normal Consistency and F-Score likewise see significant improvements. At the higher corruption level ($\sigma=0.02$), while the performance of all methods degrades, our approach remains the most robust and stable with the best mean metrics surpassing all baseline models. In challenging categories such as table, rifle, and lamp, our model still leads: rifle achieves an IoU of 0.781 and a Chamfer Distance of 0.014.
Table 1: Performance comparison of 3DS2V [48], 3DILG [47], and Ours on ShapeNet test datasets derived from the 3DS2V with an additional Gaussian noise $\sigma=0.01$. Higher is better for IoU, NC, and F-Score; lower is better for Chamfer. Best results are highlighted in bold.
@@
Category IoU $\uparrow$ NC $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ 3DS2V 3DILG Ours 3DS2V 3DILG Ours 3DS2V 3DILG Ours 3DS2V 3DILG Ours table 0.879 0.856 0.922 0.930 0.932 0.976 0.013 0.014 0.012 0.991 0.982 0.992 car 0.946 0.931 0.959 0.890 0.861 0.908 0.022 0.025 0.020 0.925 0.899 0.925 chair 0.887 0.881 0.921 0.937 0.930 0.966 0.014 0.015 0.013 0.986 0.977 0.986 airplane 0.884 0.871 0.931 0.939 0.924 0.972 0.010 0.010 0.008 0.997 0.990 0.997 sofa 0.946 0.942 0.964 0.943 0.934 0.974 0.014 0.015 0.012 0.986 0.980 0.987 rifle 0.821 0.839 0.910 0.869 0.875 0.960 0.009 0.010 0.007 0.997 0.991 0.998 lamp 0.826 0.825 0.894 0.904 0.883 0.952 0.011 0.016 0.009 0.989 0.956 0.989 mean 0.884 0.878 0.929 0.916 0.905 0.958 0.0132 0.015 0.0113 0.981 0.968 0.986
Table 2: Performance comparison of 3DS2V [48], 3DILG [47], and Ours under an additional Gaussian noise $\sigma=0.02$. Best results are highlighted in bold.
@@
Category IoU $\uparrow$ NC $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ 3DS2V 3DILG Ours 3DS2V 3DILG Ours 3DS2V 3DILG Ours 3DS2V 3DILG Ours table 0.528 0.430 0.591 0.765 0.756 0.912 0.029 0.036 0.028 0.792 0.723 0.859 car 0.434 0.541 0.491 0.715 0.699 0.787 0.040 0.052 0.044 0.669 0.600 0.688 chair 0.463 0.392 0.530 0.729 0.712 0.868 0.034 0.037 0.035 0.694 0.701 0.721 airplane 0.465 0.564 0.536 0.719 0.710 0.856 0.025 0.031 0.022 0.830 0.764 0.899 sofa 0.355 0.312 0.425 0.769 0.737 0.866 0.036 0.039 0.038 0.667 0.664 0.677 rifle 0.625 0.564 0.781 0.691 0.691 0.891 0.021 0.029 0.014 0.887 0.776 0.968 lamp 0.572 0.461 0.649 0.744 0.712 0.896 0.026 0.040 0.025 0.825 0.699 0.880 mean 0.492 0.466 0.572 0.733 0.717 0.868 0.030 0.038 0.029 0.766 0.704 0.813
@@
Figure 3: Comparison on the ShapeNet dataset. The first row corresponds to Gaussian noise with standard deviation $\sigma=0.01$, and the second row to $\sigma=0.02$. Compared to the baseline 3DS2V[48] method, our approach produces smoother reconstructions that more closely align with the underlying surfaces.
@@ 5.4 Results on ABC, Famous, and Real
We compared results on the ABC, Famous, and Real test datasets offered by P2S [15]. We compared three data-driven methods P2S [15], PPSurf [16], and 3DS2V [48] and two overfitting methods SAP-O [37], PGR [27]. These methods are widely recognized for their strong resilience to noise in point cloud data. For the data-driven methods, we used the pretrained models provided by the original authors. For the overfitting baselines, we adopted the training configurations recommended or set as default in their respective works. Quantitative results are reported in Table 3 and Table 4, and qualitative mesh reconstructions are visualized in Figure 4.
For data-driven comparison, across all noise levels our method achieves the highest NC and Mesh NC, indicating the most coherent geometry and the smoothest surfaces; this is also evident in the visual reconstructions (Figure 4). At the 0.01 noise level, the ABC and Famous datasets reach NC scores of 0.865 and 0.831, respectively—an improvement of roughly 1.5–2.5 percentage points over the second-best approach. When the noise level increases to 0.02, all baselines degrade, yet our NC remains the best among them. For methods that fit a surface to each test cloud individually, Table 4 shows that SAP-O and PGR can sometimes obtain lower Chamfer distance and higher F1, but our approach still leads on NC and Mesh NC. Under 0.01 noise level, our average NC/ Mesh NC / F-Score reach 0.847/0.027/0.945, all of which are the top scores. However, when the noise level rises to 0.02, our Chamfer and F1 decrease substantially. This drop is partly due to the backbone 3DS2V model’s sensitivity to heavy noise: its F1 plunges from 0.962 to 0.813 as the noise level increases from 0.01 to 0.02. As illustrated in Figure 4, the mesh geometry becomes severely corrupted and large holes appear. Although NoiseSDF2NoiseSDF averages the neural SDF and closes some gaps, it cannot perfectly recover the underlying surface. Moreover, the smoothing effect of our model at the 0.02 noise level removes certain fine details; while this yields visually smoother meshes, Chamfer distance and F1—metrics that emphasize global geometric fidelity rather than surface roughness—are consequently worse.
Table 3: Comparison of P2S [15], PPSurf [16], 3DS2V [48], and Ours on six noisy test datasets. Higher is better for NC, F-Score; lower is better for Mesh NC and Chamfer distance.
@@
Dataset NC $\uparrow$ Mesh NC $\downarrow$ Chamfer F-Score P2S PPSurf 3DS2V Ours P2S PPSurf 3DS2V Ours P2S PPSurf 3DS2V Ours P2S PPSurf 3DS2V Ours ABC $(\sigma=0.01)$ 0.790 0.770 0.859 0.865 0.330 0.059 0.036 0.024 0.017 0.017 0.014 0.015 0.919 0.935 0.959 0.938 ABC $(\sigma=0.02)$ 0.753 0.728 0.735 0.812 0.381 0.061 0.060 0.018 0.027 0.022 0.028 0.032 0.852 0.870 0.780 0.724 Famous $(\sigma=0.01)$ 0.771 0.761 0.819 0.831 0.268 0.053 0.040 0.025 0.017 0.015 0.014 0.016 0.928 0.959 0.962 0.941 Famous $(\sigma=0.02)$ 0.727 0.728 0.705 0.767 0.328 0.054 0.064 0.024 0.022 0.020 0.028 0.032 0.868 0.899 0.788 0.726 Real $(\sigma=0.01)$ 0.789 0.776 0.818 0.845 0.177 0.057 0.056 0.031 0.016 0.016 0.014 0.015 0.946 0.954 0.964 0.956 Real $(\sigma=0.02)$ 0.734 0.745 0.735 0.793 0.269 0.053 0.071 0.020 0.021 0.022 0.021 0.026 0.877 0.876 0.873 0.809 mean ($\sigma=0.01$) 0.783 0.769 0.832 0.847 0.258 0.056 0.044 0.027 0.017 0.016 0.014 0.015 0.931 0.949 0.962 0.945 mean (all) 0.761 0.751 0.778 0.819 0.292 0.056 0.055 0.024 0.020 0.019 0.020 0.023 0.898 0.916 0.888 0.849
Table 4: Comparison of SAP-[37], PGR [27], and Ours on six noisy test datasets. The released PGR implementation uses an adaptive Marching Cubes resolution that, for 2,048-point-point clouds, occasionally drops to 64 rather than 128, which can yield artificially smoother meshes.
@@
Dataset NC $\uparrow$ Mesh NC $\downarrow$ Chamfer F-Score SAP-O PGR Ours SAP-O PGR Ours SAP-O PGR Ours SAP-O PGR Ours ABC $(\sigma=0.01)$ 0.710 0.835 0.865 0.079 0.037 0.024 0.021 0.020 0.014 0.906 0.896 0.938 ABC $(\sigma=0.02)$ 0.622 0.778 0.812 0.095 0.065 0.018 0.026 0.026 0.032 0.824 0.815 0.724 Famous $(\sigma=0.01)$ 0.745 0.813 0.831 0.053 0.035 0.025 0.022 0.017 0.016 0.876 0.931 0.941 Famous $(\sigma=0.02)$ 0.614 0.755 0.767 0.104 0.064 0.024 0.023 0.024 0.032 0.849 0.834 0.726 Real $(\sigma=0.01)$ 0.683 0.827 0.845 0.097 0.032 0.031 0.023 0.015 0.015 0.902 0.956 0.956 Real $(\sigma=0.02)$ 0.595 0.756 0.793 0.122 0.062 0.020 0.025 0.026 0.026 0.841 0.824 0.809 mean ($\sigma=0.01$) 0.713 0.825 0.847 0.076 0.035 0.027 0.022 0.017 0.015 0.895 0.928 0.945 mean (all) 0.661 0.794 0.819 0.092 0.049 0.024 0.023 0.021 0.023 0.866 0.876 0.849
@@ 5.5 Ablation Study
Our ablation experiments were carried out on the “Chair” subset of ShapeNet, which contains 6271 models for training, 169 for validation, and 338 for testing. The configurations for all of our principal experiments originate with this subset and are progressively generalized to broader settings.
Denoising Network. The 3DS2V decoder contains 24 self-attention blocks, one cross-attention block, and a fully connected layer. Fine-tuning the entire decoder yields the best denoising performance, with metrics steadily improving over 15 epochs: IoU increases from 0.88 to 0.93, and Normal Consistency (NC) improves from 0.93 to 0.96. In comparison, fine-tuning only the fully connected layer offers virtually no benefit. Adding the cross-attention block introduces a clear gain—its final NC reaches 0.95, close to the best model—but at the cost of doubling the training time.
Table 5: Comparison of fine-tuning strategies on denoising performance.
Metric FC Layer Only FC + Cross-Attention Block Entire Decoder IoU 0.88 (no gain) 0.92 0.93 NC 0.93 (no gain) 0.95 0.96 Epochs 30 15
@@
Noise Type. Beyond standard zero-mean Gaussian noise, we evaluated three additional noise types—Uniform, Discrete, and Laplace noise—each applied at a fixed magnitude of $\sigma=0.01$. Furthermore, to assess the impact of non-zero bias in Gaussian perturbations, we conducted experiments over the domain $[-1,1]^{3}$ using means of $\mu=0.005$, $0.01$, and $0.02$. Comprehensive quantitative results are presented in Table 6.
At $\sigma=0.01$ on $[-1,1]^{3}$, our model consistently shows denoising performance under Uniform and Discrete noise, with notable gains in both IoU and NC. However, it shows limited benefit under Laplace noise. For Gaussian noise with $\sigma=0.01$ increasing $\mu$, our model remains effective at lower $\mu$, but its advantage diminishes as the $\mu$ grows, eventually leading to degraded reconstruction quality.
Table 6: Performance comparison between the 3DS2V [48] and our method under various noise types and levels. Noise distributions tested are Uniform, Discrete and Laplace with $\sigma=0.01$, and Gaussian with $\sigma=0.01$ under different means $\mu=0,0.005,0.01,0.02$.
@@
Dataset IoU $\uparrow$ NC $\uparrow$ Chamfer F-Score 3DS2V Ours 3DS2V Ours 3DS2V Ours 3DS2V Ours Uniform $(\sigma=0.01)$ 0.873 0.911 0.920 0.962 0.015 0.013 0.985 0.986 Discrete $(\sigma=0.01)$ 0.867 0.895 0.915 0.960 0.015 0.014 0.984 0.985 Laplace $(\sigma=0.01)$ 0.909 0.908 0.956 0.956 0.014 0.014 0.985 0.985 Gaussian $(\sigma=0.01)$ 0.887 0.927 0.937 0.966 0.014 0.013 0.986 0.986 Gaussian $(\sigma=0.01,\mu=0.005)$ 0.860 0.881 0.942 0.964 0.017 0.016 0.982 0.984 Gaussian $(\sigma=0.01,\mu=0.01)$ 0.798 0.800 0.946 0.949 0.024 0.023 0.952 0.960 Gaussian $(\sigma=0.01,\mu=0.02)$ 0.661 0.666 0.875 0.898 0.038 0.039 0.549 0.524

### v2

@@ 5.3 Results on ShapeNet
To verify our hypothesis that clean neural fields can be learned from noisy supervision, we comprehensively compare our method with 3DS2V [50] on the seven largest ShapeNet [8] subsets, following its experimental setup. Since our method is trained using 3DS2V as the denoising network with only noisy supervision, achieving superior performance over 3DS2V would therefore provide strong evidence supporting our hypothesis.
We adopt the officially released 3DS2V model as our baseline. 3DS2V was trained with pairs of clean point clouds and clean SDFs for supervision, and it was not exposed to noisy inputs paired with clean SDF. To ensure a fair evaluation, we adopt the same data splits and preprocessing procedures as 3DS2V. Our NoiseSDF2NoiseSDF is trained under noisy supervision, meaning that we do not use any paired noisy point clouds and clean SDFs throughout the entire training process.
We report evaluation results for each subset at noise levels of 0.01 (Table 1) and 0.02 (Table 2) and show visualization results in Figure 3. Under lower corruption ($\sigma=0.01$), our method outperforms 3DS2V across all evaluation metrics. For example, the mean results show an IoU increase of 0.045 (5.1%), a Normal Consistency improvement of 0.042 (4.6%), and a reduction in Chamfer Distance from 0.0132 to 0.0113. At the higher corruption level ($\sigma=0.02$), our approach remains the robust and stable with the better mean metrics that surpass the baseline. For instance, the mean results show an IoU increase of 0.08 (16.3%), a Normal Consistency improvement of 0.135 (18.4%), and an F-Score increase of 0.047 (6%). The results demonstrate that our method outperforms the baseline, 3DS2V, under both corruption levels and confirm our central idea that it is possible to learn to clean neural fields from noisy supervision.
Table 1: Comparison of 3DS2V [50] and Ours on ShapeNet with Gaussian noise $\sigma=0.01$.
@@
Category IoU $\uparrow$ NC $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ table 0.879 0.922 +0.043 0.930 0.976 +0.046 0.013 0.012 +0.001 0.991 0.992 +0.001 car 0.946 0.959 +0.013 0.890 0.908 +0.018 0.022 0.020 +0.002 0.925 0.925 +0.000 chair 0.887 0.921 +0.034 0.937 0.966 +0.029 0.014 0.013 +0.001 0.986 0.986 +0.000 airplane 0.884 0.931 +0.047 0.939 0.972 +0.033 0.010 0.008 +0.002 0.997 0.997 +0.000 sofa 0.946 0.964 +0.018 0.943 0.974 +0.031 0.014 0.012 +0.002 0.986 0.987 +0.001 rifle 0.821 0.910 +0.089 0.869 0.960 +0.091 0.009 0.007 +0.002 0.997 0.998 +0.001 lamp 0.826 0.894 +0.068 0.904 0.952 +0.048 0.011 0.009 +0.002 0.989 0.989 +0.000 mean 0.884 0.929 +0.045 0.916 0.958 +0.042 0.0132 0.0113 +0.0019 0.981 0.986 +0.005
Table 2: Comparison of 3DS2V [50] and Ours on ShapeNet with Gaussian noise $\sigma=0.02$.
@@
Category IoU $\uparrow$ NC $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ table 0.528 0.591 +0.063 0.765 0.912 +0.147 0.029 0.028 +0.001 0.792 0.859 +0.067 car 0.434 0.491 +0.057 0.715 0.787 +0.072 0.040 0.044 -0.004 0.669 0.688 +0.019 chair 0.463 0.530 +0.067 0.729 0.868 +0.139 0.034 0.035 -0.001 0.694 0.721 +0.027 airplane 0.465 0.536 +0.071 0.719 0.856 +0.137 0.025 0.022 +0.003 0.830 0.899 +0.069 sofa 0.355 0.425 +0.070 0.769 0.866 +0.097 0.036 0.038 -0.002 0.667 0.677 +0.010 rifle 0.625 0.781 +0.156 0.691 0.891 +0.200 0.021 0.014 +0.007 0.887 0.968 +0.081 lamp 0.572 0.649 +0.077 0.744 0.896 +0.152 0.026 0.025 +0.001 0.825 0.880 +0.055 mean 0.492 0.572 +0.080 0.733 0.868 +0.135 0.030 0.029 +0.001 0.766 0.813 +0.047
@@
Figure 3: Comparison on the ShapeNet dataset. The first row shows Gaussian noise with $\sigma=0.01$, and the second row with $\sigma=0.02$. Ours produces smoother reconstructions, better aligning with the underlying surfaces compared to baseline.
@@ 5.4 Results on ABC, Famous, and Real
We compared results on the ABC, Famous, and Real test datasets provided by P2S [15]. Specifically, we evaluated data-driven methods P2S, PPSurf [16], POCO [6], and NKSR [22], known for their strong noise resilience in point cloud data. Note that, except for our method, all other approaches are trained on the ABC dataset directly using ground-truth noisy–clean pairs. For these methods, we used their officially released pretrained models. Quantitative results are reported in Table 3, and qualitative mesh reconstructions are shown in Figure 4.
Across noise levels, our method achieves strong performance on mean NC and Mesh NC, indicating coherent geometry and smooth surfaces; this is also evident in the visual reconstructions (Figure 4). At the 0.01 noise level, the ABC and Famous datasets achieve NC scores of 0.865 and 0.831, respectively. When the noise level increases to 0.02, our method yields the best mean Mesh NC and strong mean NC among them. These results demonstrate that even when trained with noisy supervision and without ground-truth, our method can achieve performance competitive with state-of-the-art data-driven approaches.
5.5 Quantitative Results Supplementary
Table 3: Comparison of P2S [15], PPSurf [16], POCO [6], NKSR [22], and Ours on six noisy test datasets. The best two performances are highlighted.
@@ 5.5 Quantitative Results Supplementary
Dataset NC $\uparrow$ Mesh NC $\downarrow$ Chamfer $\downarrow$ F-Score $\uparrow$ P2S PPSurf POCO NKSR Ours P2S PPSurf POCO NKSR Ours P2S PPSurf POCO NKSR Ours P2S PPSurf POCO NKSR Ours ABC $(\sigma=0.01)$ 0.790 0.770 0.864 0.800 0.865 0.330 0.059 0.025 0.039 0.024 0.017 0.017 0.014 0.018 0.015 0.919 0.935 0.941 0.927 0.938 ABC $(\sigma=0.02)$ 0.753 0.728 0.848 0.727 0.812 0.381 0.061 0.020 0.058 0.018 0.027 0.022 0.019 0.026 0.032 0.852 0.870 0.898 0.801 0.724 Famous $(\sigma=0.01)$ 0.771 0.761 0.825 0.775 0.831 0.268 0.053 0.026 0.032 0.025 0.017 0.015 0.017 0.017 0.016 0.928 0.959 0.962 0.943 0.941 Famous $(\sigma=0.02)$ 0.727 0.728 0.785 0.703 0.767 0.328 0.054 0.023 0.054 0.024 0.022 0.020 0.022 0.026 0.032 0.868 0.899 0.871 0.810 0.726 Real $(\sigma=0.01)$ 0.789 0.776 0.845 0.779 0.845 0.177 0.057 0.032 0.038 0.031 0.016 0.016 0.014 0.015 0.015 0.946 0.954 0.964 0.955 0.956 Real $(\sigma=0.02)$ 0.734 0.745 0.803 0.700 0.793 0.269 0.053 0.024 0.059 0.020 0.021 0.022 0.025 0.029 0.026 0.877 0.876 0.930 0.822 0.809 mean ($\sigma=0.01$) 0.783 0.769 0.844 0.785 0.847 0.258 0.056 0.028 0.036 0.027 0.017 0.014 0.016 0.017 0.015 0.931 0.949 0.962 0.930 0.945 mean (all) 0.761 0.751 0.828 0.747 0.819 0.292 0.056 0.025 0.047 0.024 0.020 0.019 0.019 0.022 0.023 0.898 0.916 0.915 0.876 0.849
We compare our method with representative overfitting-based approaches, such as SAP-O [39] and PGR [29]. These methods train separate networks for each test shape, require long inference times, and lack generalization to unseen shapes. We adopted the training configurations recommended or set as default in their respective works. Table 4 shows that SAP-O and PGR can sometimes achieve lower Chamfer Distance and higher F1, but our approach consistently outperforms them in NC and Mesh NC. At the 0.01 noise level, our method achieves average NC/Mesh NC/F-Score values of 0.847/0.027/0.945, all of which are the best scores. At the 0.02 noise level, our method also demonstrates strong performance, with higher NC and Mesh NC, as well as competitive Chamfer Distance and F-Score.
Table 4: Comparison of SAP‑O [39], PGR [29], and Ours on six noisy test datasets. The adaptive resolution in PGR can occasionally drop relatively low, resulting in artificially smooth meshes.
@@
Dataset NC $\uparrow$ Mesh NC $\downarrow$ Chamfer $\downarrow$ F‑Score $\uparrow$ SAP‑O PGR Ours SAP‑O PGR Ours SAP‑O PGR Ours SAP‑O PGR Ours ABC $(\sigma=0.01)$ 0.710 0.835 0.865 0.079 0.037 0.024 0.021 0.020 0.014 0.906 0.896 0.938 ABC $(\sigma=0.02)$ 0.622 0.778 0.812 0.095 0.065 0.018 0.026 0.026 0.032 0.824 0.815 0.724 Famous $(\sigma=0.01)$ 0.745 0.813 0.831 0.053 0.035 0.025 0.022 0.017 0.016 0.876 0.931 0.941 Famous $(\sigma=0.02)$ 0.614 0.755 0.767 0.104 0.064 0.024 0.023 0.024 0.032 0.849 0.834 0.726 Real $(\sigma=0.01)$ 0.683 0.827 0.845 0.097 0.032 0.031 0.023 0.015 0.015 0.902 0.956 0.956 Real $(\sigma=0.02)$ 0.595 0.756 0.793 0.122 0.062 0.020 0.025 0.026 0.026 0.841 0.824 0.809 mean ($\sigma=0.01$) 0.713 0.825 0.847 0.076 0.035 0.027 0.022 0.017 0.015 0.895 0.928 0.945 mean (all) 0.661 0.794 0.819 0.092 0.049 0.024 0.023 0.021 0.023 0.866 0.876 0.849
Table 5: Comparison of N2NM and our method under medium noise (MedN) and maximum noise (MaxN) using Chamfer Distance on the Famous dataset.
Method MedN Chamfer $\downarrow$ MaxN Chamfer $\downarrow$ Inference time
N2NM 0.0132 0.0231 2760 seconds
Ours 0.0160 0.0320 0.05 seconds
Ours-TTO 0.0108 0.0252 20 seconds
We further include a comparison with N2NM [33]. For each shape, N2NM was trained with 200 noisy samples, which resulted in long inference times (46 minutes) per shape. To ensure a fair comparison, we conducted experiments on the Famous dataset following the N2NM setup [53] and incorporated test-time optimization (TTO) into our method with the same number of noisy samples. As shown in Table 5, our method combined with TTO demonstrated significant self-improvement, outperformed N2NM under medium noise conditions, and achieved comparable results under maximum noise, while delivering inference that is orders of magnitude faster.
@@ 5.6 Ablation Study
Our ablation experiments use the “Chair” subset of ShapeNet, with 6271 models for training, 169 for validation, and 338 for testing.
Noisy vs. Clean Supervision. In the context of Noise2Noise, clean supervision corresponds to Noise2Clean, where ground-truth SDFs are provided during training. For a fair comparison, we conducted experiments under identical training conditions (batch size, shapes numbers/batch, etc.) between our noisy and clean supervision with a noise level of 0.01. The results are reported in Table 6(b) (Clean as target). We observe that our noisy supervision achieves performance nearly equivalent to that of clean supervision, which aligns with the findings of Lehtinen et al. [27]. This supports our main hypothesis that cleaner outputs can be produced by observing noisy neural fields.
Noisy Targets and Denoising Networks. To validate that our framework can relax constraints on the architecture of the SDF prediction (i.e., Point2SDF), we conducted a study by replacing 3DS2V with 3DILG as the noisy targets. We further used 3DILG as the denoising network and performed same experiments. 3DILG [49] encodes 3D shapes using irregular latent grids, whereas 3DS2V represents neural fields with a set of vectors. In Table 6(b), the results indicate that our method remains effective in learning clean SDFs from noisy supervision. These findings support the generalizability of our approach across different forms of noisy supervision and denoising networks.
Table 6: Quantitative comparisons for different targets and denoising networks.
(a) Using 3DS2V as denoising network.
Method IoU $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ NC $\uparrow$
Baseline 0.887 0.014 0.986 0.937
Clean 0.939 0.013 0.988 0.970
Noisy(3DILG) 0.903 0.014 0.980 0.950
Noisy(3DS2V) 0.927 0.013 0.986 0.966
(b) Using 3DILG as denoising network.
Method IoU $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ NC $\uparrow$
Baseline 0.881 0.015 0.977 0.930
Clean 0.921 0.014 0.981 0.960
Noisy(3DILG) 0.907 0.014 0.979 0.953
Noisy(3DS2V) 0.913 0.014 0.978 0.962
Denoising Network Training. Fine-tuning only the fully connected layer offers virtually no benefit. Adding the cross-attention block introduces a clear gain. Fine-tuning the entire decoder achieves the better performance. We further evaluate the random initialization of the decoder and the entire network. Table 7 shows that initializing the decoder from 3DS2V does not improve performance but does accelerate convergence. Considering the trade-off between performance and training cost, we adopt the strategy of freezing the encoder while fine-tuning the decoder.
Table 7: Comparison of training strategies on the denoising network.
Metric Baseline FC FC+CA Decoder Decoder-RI Network-RI IoU 0.887 0.884 0.922 0.927 0.927 0.929 NC 0.937 0.933 0.954 0.966 0.967 0.968 Epochs 0 10 15 30 350 800
@@
Noise Type. Beyond standard zero‐mean Gaussian noise, we evaluated three additional noise types, Uniform, Discrete, and Laplace noise, each applied at a fixed magnitude of $\sigma=0.01$. Furthermore, to assess the impact of non-zero bias in Gaussian perturbations, we conducted experiments over the domain $1,1]^{3}$ using means of $\mu=0.005$, $0.01$, and $0.02$. Comprehensive quantitative results are presented in Table 8. At $\sigma=0.01$ on $1,1]^{3}$, our model consistently shows denoising performance under Uniform and Discrete noise, with notable gains in both IoU and NC.
Table 8: Comparison between the 3DS2V [50] and Ours under various noise types.
@@
Dataset IoU $\uparrow$ NC $\uparrow$ Chamfer $\downarrow$ F-Score $\uparrow$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ 3DS2V Ours $\Delta$ Uniform $(\sigma=0.01)$ 0.873 0.911 +0.038 0.920 0.962 +0.042 0.015 0.013 +0.002 0.985 0.986 +0.001 Discrete $(\sigma=0.01)$ 0.867 0.895 +0.028 0.915 0.960 +0.045 0.015 0.014 +0.001 0.984 0.985 +0.001 Laplace $(\sigma=0.01)$ 0.909 0.908 -0.001 0.956 0.956 +0.000 0.014 0.014 +0.000 0.985 0.985 +0.000 Gaussian $(\sigma=0.01)$ 0.887 0.927 +0.040 0.937 0.966 +0.029 0.014 0.013 +0.001 0.986 0.986 +0.000 Gaussian $(\sigma=0.01,\mu=0.005)$ 0.860 0.881 +0.021 0.942 0.964 +0.022 0.017 0.016 +0.001 0.982 0.984 +0.002 Gaussian $(\sigma=0.01,\mu=0.01)$ 0.798 0.800 +0.002 0.946 0.949 +0.003 0.024 0.023 +0.001 0.952 0.960 +0.008 Gaussian $(\sigma=0.01,\mu=0.02)$ 0.661 0.666 +0.005 0.875 0.898 +0.023 0.038 0.039 -0.001 0.549 0.524 -0.025 mean 0.836 0.855 +0.019 0.927 0.950 +0.023 0.020 0.019 +0.001 0.917 0.912 -0.005